OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (6, 3) and has a slope of ? a. 2x + 3y = 3 b. 2x – 3y = 3 c. 3x + 2y = 3 d. 3x – 2y = 3

OpenStudy (anonymous):

The equation for a line is: \[y-y_1=m(x-x_1)\]Where m=slope, x_1,y_1 are the coordinates of a point on the line. Plug in your values and rearrange as necessary.